# Structured Probabilistic Inference & Generative Modeling

### v1

@@ 1 Introduction
Over the last few years there has been tremendous growth at the intersection of scientific machine learning (SciML), traditional scientific computing methods, and computer science algorithms. In particular, there has been growing interest in utilizing neural networks to learn physical operators of dynamical systems, such as high-dimensional partial differential equations (PDEs) Lu et al. [2022a], Kovachki et al. [2023]. These methods are referred to under the general category of operator learning. However, in order for these operator learning approaches to achieve significant impact in scientific and engineering applications, they must have reliable and robust methods for assessing the uncertainties associated with training and deploying models.
@@
Uncertainties from machine learning models are often placed into two categories: aleatoric uncertainties arising from inherent stochasticities, e.g., in stochastic ODEs and PDEs; and epistemic uncertainties arising from lack of data. Ref. Psaros et al. [2023] provides a thorough review of the types of uncertainties that can influence solutions to various dynamical systems obtained via SciML methods, such as operator learning. In this paper, we focus on the case of aleatoric uncertainties. In particular, we are interested in learning an operator that accurately estimates predictive uncertainty due to such variability. That is, given a fixed input $x$, what is the probability distribution $p$ that best describes the action of a stochastic operator $\mathcal{M}$ acting on $x$?
@@
Many current approaches, including Bayesian and Frequentist frameworks, rely on a likelihood function to model aleatoric uncertainty. The stochastic operator $\mathcal{M}$ is parameterized by some model architecture. Optionally, a prior probability distribution can also be placed on the parameters of the model. By assuming a form for the likelihood function on the data-model relationship, the parameters of the model can be found via maximum likelihood estimation (MLE) or by maximum a posteriori (MAP) estimation Psaros et al. [2023]. Often, the likelihood is also parameterized and optimized to more flexibly model the stochastic behavior of the system Kendall and Gal [2017]. These sample parameters can then be utilized to evaluate the model and, when convolved with the likelihood function, be used to estimate the predictive uncertainty of the output of the operator. However, these approaches can be computationally prohibitive to execute, and approximation methods are often used, such as assuming no correlation between model outputs Kendall and Gal [2017]. In addition, the computational complexity of likelihood based approaches, they often also require strong assumptions about the likelihood function which may be inappropriate for assessing the type of uncertainties relevant to systems of stochastic ODEs and PDEs.
@@
Recently, generative modeling has demonstrated success across a variety of machine learning tasks Bengesi et al. [2024]. These models can also be interpreted as providing uncertainty quantification Mimikos-Stamatopoulos et al. [2024]. Score-based approaches towards generative modeling are particularly attractive because they neither require a likelihood nor involve adversarial optimization Pacchiardi and Dutta [2022]. Authors have used generative models as hyper-networks to generate the weights for a neural network model, allowing one to estimate uncertainties for model predictions Krueger et al. [2017], Ratzlaff and Fuxin [2019], Safta et al. [2025]. These works, however, did not provide isolated analysis of quantified aleatoric uncertainties apart from epistemic uncertainties. Moreover, instead of using score based approaches to generative modeling, they relied either on generative adversarial approaches or methods that require likelihood functions. Finally, none of these works have applied these hyper-network UQ approaches towards operator learning.
@@ 2 GenUQ: Predictive Uncertainty from Generative Models
Similar to Bayesian frameworks such as Neal [2012], Meng et al. [2021], we assume that the parameters $\theta$ are random variables and can model the dataset by providing a probability measure for $\mathbb{P}_{\Theta}$ or a means of sampling $\theta\sim\mathbb{P}_{\Theta}$ such that its pushforward matches the data. However, instead of following the traditional frameworks of finding a distribution for $\theta$ given the data through Bayes’ Rule, we instead seek to estimate a measure $\mathbb{P}_{\Theta}$ whose push-forward approximates the output target measure $\mathbb{P}_{\mathcal{W}}(W)$. That is,
@@
where $h^{-1}(W)=\{\theta\in\Theta:h_{U}(\theta)=(U,f_{\theta}(U))=W\}$. This approach bears similarity to other measure-theoretic frameworks for uncertainty quantification Butler et al. [2018, 2025].
@@
As discussed in the following section, there are many options for learning such a sampling distribution, including normalizing flows Rezende and Mohamed [2015a]. In our experiments, we find that utilizing a non-invertible feed-forward neural network for $g_{\phi}$ is sufficient provided an appropriate discrepancy function. To maintain the computational efficiency of GenUQ, we seek a generative model for only a small, random subset of the underlying model’s parameters, leaving the remainder as deterministic variables. The proportion of generated parameters, $R$, is a hyperparameter that we examine in the results section.
@@
where $\beta\in(0,2)$. Note that the two terms in the norm, $\|\cdot\|_{\mathcal{W}}$ can be weighted if $\|u\|_{\mathcal{U}}/\|v\|_{\mathcal{V}}\neq\mathcal{O}(1)$. This score is associated with the maximum-mean discrepancy (MMD) metric for a specific choice of kernel distance Bińkowski et al. [2018]. There are a number of advantages to utilizing such a discrepancy score to train models. See Gneiting and Raftery [2007], Pacchiardi and Dutta [2022] for details. In particular, one of the advantages of using this type of discrepancy score is that explicit knowledge of the underlying distributions is not required to approximate the energy score (as with other MMD metrics), i.e., the energy score can be approximated via Monte Carlo from finite samples of the two comparison distributions $\mathbb{P}$ and $\mathbb{Q}$. Indeed, we have samples from the joint distribution $\{w_{i}=(u_{i},v_{i})\}\sim\mathbb{P}_{\mathcal{W}}$ in practice. Thus, we can train the hyper-network generative model with the following criteria:
@@ 3 Related Works
Bayesian approaches for uncertainty quantification are very popular, especially in the scientific machine learning community. Among them, variational inference is commonly chosen because it enables inference for posterior parameters in situations where performing MCMC sampling is challenging Graves [2011], Blei et al. [2017]. Nonetheless, classical variational inference approaches can be limited by assumptions on the class of approximation densities used to approximate the posterior (e.g., mean-field approximations). To overcome this limitation, strategies such as normalizing flows have been proposed to generate a more flexible class of approximation densities for use in the variational inference approach Rezende and Mohamed [2015b]. Our approach is in a similar spirit to this type of approach but differs in the mathematical framework underlying the approach. While variational inference also seeks a distribution over model parameters, the posterior distribution it attempts to approximate is assumed to be a convolution of the data likelihood and prior distribution. Instead, we seek a distribution over the model parameters whose push-forward measure is consistent with the distribution of the data, i.e., via the discrepancy score (4). These differences in objectives can lead to significantly different mathematical consequences–see Butler et al. [2018] for details. Our approach is most similar to Lingsch et al. [2024], which learns a solution to an inverse problem (in addition to the forward operator) to characterize the stochastic inputs to a PDE. They associate every input $u\in\mathcal{U}$ with a specific stochastic parameterization $\xi\mid u\in\Xi$, and then learn the appropriate mapping $f_{\theta}:\Xi\rightarrow\mathcal{V}$. On the other hand, we assume that $\xi$ is independent of $u$ and directly influences the output $v$ through the operator. Thus, our method, GenUQ, focuses on a forward problem directly relating $\mathcal{U}$ and $\mathcal{V}$ (via $f_{\theta}:\mathcal{U}\rightarrow\mathcal{V}$), bypassing the need to solve an extra inverse problem.
@@ 4 Results
Throughout these examples, we compare to 5 other UQ methods. The first three are maximum likelihood estimation based methods where we assume Gaussian error with heteroscadastic, but uncorrelated noise in a mean field variational inference model (VI), heteroscadastic noise with full covariance structure (CoV), and normalizing flow conditioned on the input function (NF). While VI provides both epistemic and aleatoric uncertainties, we include it due to its ubiquity in machine learning. For CoV, we predict a full covariance matrix from the neural operator. For NF, we use masked autoregressive flow Papamakarios et al. [2017] from the error distribution to a normal distribution using the FlowJax library with default options Ward [2025].
@@ 4.1 ELU Operator learning example
We compare GenUQ to three other UQ methods in recovering a manufactured stochastic operator from its action on random smooth functions. The operator we use is the composition between the exponential linear unit (ELU) point-wise non-linearity Clevert et al. [2016] and a differentiation,
@@
Our underlying model is the operator learning method, Modal Operator Regression for Physics (MOR-Physics)Patel and Desjardins [2018], Patel et al. [2021], ${f}_{\theta}(u)=\mathcal{F}^{-1}g_{\theta_{g}}(\kappa)\mathcal{F}h_{\theta_{h}}(u)$, where $\mathcal{F}$ is the Fourier transform, $\kappa$ is the wavenumber, and $g$ and $h$ are neural networks. For both neural networks, we use a depth of 6 and width of 32 and the GELU activation function. For the generator network, we use a four layer network of widths, $(10,20,30,40)$, and the same activation function.
@@ 4.2 Stochastic Poisson equation solution operator
We will refer to the stochastic solution operator for (6) provided by the finite element solver as $N_{true}$ and seek a neural operator, $N$, that emulates it. We parameterize the neural operator with the POD-DeepONets architecture with hard constrained boundary conditions Lu et al. [2022b]. We construct a truncated POD basis of dimension $d=200$ from the training input functions, and project the input functions onto this basis to produce the POD coefficients, $c$. The POD-DeepONets architecture has the following form,
@@
where $t$ and $b$ are trunk and branch neural networks of depth 5 and width 128. See Lu et al. [2022b] for further details on this architecture.

### v2

@@ 1 Introduction
Over the last few years there has been tremendous growth at the intersection of scientific machine learning (SciML), traditional scientific computing methods, and computer science algorithms. In particular, there has been growing interest in utilizing neural networks to learn physical operators of dynamical systems, such as high-dimensional partial differential equations (PDEs) patel2018nonlinear, patel2021morp, LMC22-OL, KLL23-OL. These methods are referred to under the general category of operator learning. However, in order for these operator learning approaches to achieve significant impact in scientific and engineering applications, they must have reliable and robust methods for assessing the uncertainties associated with training and deploying models.
@@
Uncertainties from machine learning models are often placed into two categories: aleatoric uncertainties arising from inherent stochasticities, e.g., in stochastic ODEs and PDEs; and epistemic uncertainties arising from lack of data. Ref. PMZ23-UQSciML provides a thorough review of the types of uncertainties that can influence solutions to various dynamical systems obtained via SciML methods, such as operator learning. In this paper, we focus on the case of aleatoric uncertainties. In particular, we are interested in learning an operator that accurately estimates predictive uncertainty due to such variability. That is, given a fixed input $x$, what is the probability distribution $p$ that best describes the action of a stochastic operator $\mathcal{M}$ acting on $x$?
@@
Many current approaches, including Bayesian and Frequentist frameworks, rely on a likelihood function to model aleatoric uncertainty. The stochastic operator $\mathcal{M}$ is parameterized by some model architecture. Optionally, a prior probability distribution can also be placed on the parameters of the model. By assuming a form for the likelihood function on the data-model relationship, the parameters of the model can be found via maximum likelihood estimation (MLE) or by maximum a posteriori (MAP) estimation PMZ23-UQSciML. Often, the likelihood is also parameterized and optimized to more flexibly model the stochastic behavior of the system kendalluq. These sample parameters can then be utilized to evaluate the model and, when convolved with the likelihood function, be used to estimate the predictive uncertainty of the output of the operator. However, these approaches can be computationally prohibitive to execute, and approximation methods are often used, such as assuming no correlation between model outputs kendalluq. In addition, the computational complexity of likelihood based approaches, they often also require strong assumptions about the likelihood function which may be inappropriate for assessing the type of uncertainties relevant to systems of stochastic ODEs and PDEs.
@@
Recently, generative modeling has demonstrated success across a variety of machine learning tasks bengesi2024gm. These models can also be interpreted as providing uncertainty quantification mimikos-stamatopoulos2024scorebased. Score-based approaches towards generative modeling are particularly attractive because they neither require a likelihood nor involve adversarial optimization pacchiardi2022score. Authors have used generative models as hyper-networks to generate the weights for a neural network model, allowing one to estimate uncertainties for model predictions krueger2017bayesian, ratzlaff2019hypergan, safta2025uncertainty. These works, however, did not provide isolated analysis of quantified aleatoric uncertainties apart from epistemic uncertainties. Moreover, instead of using score based approaches to generative modeling, they relied either on generative adversarial approaches or methods that require likelihood functions. Finally, none of these works have applied these hyper-network UQ approaches towards operator learning.
@@ 2 GenUQ: Predictive Uncertainty from Generative Models
Similar to Bayesian frameworks such as N12-BNNThesis, MBK21-MFBNN, we assume that the parameters $\theta$ are random variables and can model the dataset by providing a probability measure for $\mathbb{P}_{\Theta}$ or a means of sampling $\theta\sim\mathbb{P}_{\Theta}$ such that its pushforward matches the data. However, instead of following the traditional frameworks of finding a distribution for $\theta$ given the data through Bayes’ Rule, we instead seek to estimate a measure $\mathbb{P}_{\Theta}$ whose push-forward approximates the output target measure $\mathbb{P}_{\mathcal{W}}(W)$. That is,
@@
where $h^{-1}(W)=\{\theta\in\Theta:h_{U}(\theta)=(U,f_{\theta}(U))=W\}$. This approach bears similarity to other measure-theoretic frameworks for uncertainty quantification BJW18-DCIClassic, BSW25-DCIRecent.
@@
As discussed in the following section, there are many options for learning such a sampling distribution, including normalizing flows rezende2015variational. In our experiments, we find that utilizing a non-invertible feed-forward neural network for $g_{\phi}$ is sufficient provided an appropriate discrepancy function. To maintain the computational efficiency of GenUQ, we seek a generative model for only a small, random subset of the underlying model’s parameters, leaving the remainder as deterministic variables. The proportion of generated parameters, $R$, is a hyperparameter that we examine in the results section.
@@
where $\beta\in(0,2)$. Note that the two terms in the norm, $\|\cdot\|_{\mathcal{W}}$ can be weighted if $\|u\|_{\mathcal{U}}/\|v\|_{\mathcal{V}}\neq\mathcal{O}(1)$. This score is associated with the maximum-mean discrepancy (MMD) metric for a specific choice of kernel distance BDS18-MMDGans. There are a number of advantages to utilizing such a discrepancy score to train models. See gneiting2007strictly, pacchiardi2022score for details. In particular, one of the advantages of using this type of discrepancy score is that explicit knowledge of the underlying distributions is not required to approximate the energy score (as with other MMD metrics), i.e., the energy score can be approximated via Monte Carlo from finite samples of the two comparison distributions $\mathbb{P}$ and $\mathbb{Q}$. Indeed, we have samples from the joint distribution $\{w_{i}=(u_{i},v_{i})\}\sim\mathbb{P}_{\mathcal{W}}$ in practice. Thus, we can train the hyper-network generative model with the following criteria:
@@ 3 Related Works
Bayesian approaches for uncertainty quantification are very popular, especially in the scientific machine learning community. Among them, variational inference is commonly chosen because it enables inference for posterior parameters in situations where performing MCMC sampling is challenging G11-VI4NN, BKM17-VIReview. Nonetheless, classical variational inference approaches can be limited by assumptions on the class of approximation densities used to approximate the posterior (e.g., mean-field approximations). To overcome this limitation, strategies such as normalizing flows have been proposed to generate a more flexible class of approximation densities for use in the variational inference approach RM15-VIwNF. Our approach is in a similar spirit to this type of approach but differs in the mathematical framework underlying the approach. While variational inference also seeks a distribution over model parameters, the posterior distribution it attempts to approximate is assumed to be a convolution of the data likelihood and prior distribution. Instead, we seek a distribution over the model parameters whose push-forward measure is consistent with the distribution of the data, i.e., via the discrepancy score (4). These differences in objectives can lead to significantly different mathematical consequences–see BJW18-DCIClassic for details. Our approach is most similar to LGM24-fuse, which learns a solution to an inverse problem (in addition to the forward operator) to characterize the stochastic inputs to a PDE. They associate every input $u\in\mathcal{U}$ with a specific stochastic parameterization $\xi\mid u\in\Xi$, and then learn the appropriate mapping $f_{\theta}:\Xi\rightarrow\mathcal{V}$. On the other hand, we assume that $\xi$ is independent of $u$ and directly influences the output $v$ through the operator. Thus, our method, GenUQ, focuses on a forward problem directly relating $\mathcal{U}$ and $\mathcal{V}$ (via $f_{\theta}:\mathcal{U}\rightarrow\mathcal{V}$), bypassing the need to solve an extra inverse problem.
@@ 4 Results
Throughout these examples, we compare to 5 other UQ methods. The first three are maximum likelihood estimation based methods where we assume Gaussian error with heteroscadastic, but uncorrelated noise in a mean field variational inference model (VI), heteroscadastic noise with full covariance structure (CoV), and normalizing flow conditioned on the input function (NF). While VI provides both epistemic and aleatoric uncertainties, we include it due to its ubiquity in machine learning. For CoV, we predict a full covariance matrix from the neural operator. For NF, we use masked autoregressive flow papamakarios2017masked from the error distribution to a normal distribution using the FlowJax library with default options ward2023flowjax.
@@ 4.1 ELU Operator learning example
We compare GenUQ to three other UQ methods in recovering a manufactured stochastic operator from its action on random smooth functions. The operator we use is the composition between the exponential linear unit (ELU) point-wise non-linearity clevert2016elu and a differentiation,
@@
Our underlying model is the operator learning method, Modal Operator Regression for Physics (MOR-Physics)patel2018nonlinear, patel2021morp, ${f}_{\theta}(u)=\mathcal{F}^{-1}g_{\theta_{g}}(\kappa)\mathcal{F}h_{\theta_{h}}(u)$, where $\mathcal{F}$ is the Fourier transform, $\kappa$ is the wavenumber, and $g$ and $h$ are neural networks. For both neural networks, we use a depth of 6 and width of 32 and the GELU activation function. For the generator network, we use a four layer network of widths, $(10,20,30,40)$, and the same activation function.
@@ 4.2 Stochastic Poisson equation solution operator
We will refer to the stochastic solution operator for (6) provided by the finite element solver as $N_{true}$ and seek a neural operator, $N$, that emulates it. We parameterize the neural operator with the POD-DeepONets architecture with hard constrained boundary conditions Lupod-don2022. We construct a truncated POD basis of dimension $d=200$ from the training input functions, and project the input functions onto this basis to produce the POD coefficients, $c$. The POD-DeepONets architecture has the following form,
@@
where $t$ and $b$ are trunk and branch neural networks of depth 5 and width 128. See Lupod-don2022 for further details on this architecture.